OpenStudy (anonymous):

A cylindrical well is 15 meters deep and has a diameter of 1.6 meters. Approximately how many cubic meters of soil were dug out to make the well? (Use π = 3.14.) Is anyone woke i need help

OpenStudy (anonymous):

30.14 cubic meters 37.68 cubic meters 120.57 cubic meters 188.40 cubic meters

OpenStudy (baru):

you know the formula for volume of a cylinder?

OpenStudy (er.mohd.amir):

volume of a cylinder=pi*r*r*h

OpenStudy (er.mohd.amir):

use this and cal.

OpenStudy (er.mohd.amir):

where r is radius and h is height and radius is half of daimeter

OpenStudy (anonymous):

Sorry i didn't know someone reply back, i'm not good at these type of questions

OpenStudy (er.mohd.amir):

r=d/2=1.6/2=0.8 volume of a cylinder=pi*r*r*h =3.14*0.8*0.8*15=?

OpenStudy (anonymous):

30.144 that's the answer i got

OpenStudy (er.mohd.amir):

right

OpenStudy (anonymous):

Thank you soo much